A solution was made by dissolving 10.0 g of potassium sulfate in 250.00 mL of solution.

1. How many moles of potassium sulfate are in solution?

2. How many moles of each ion are in solution?

3. What is the concentration of potassium sulfate? Each ion?
Answer #1

1.weight of K2SO4   = 10g

gram molecular weight of K2SO4 = 174.259g/mole

no of moles of K2SO4   = .weight of K2SO4/gram molecular weight of K2SO4

                                       = 10/174.259    = 0.0574moles

2.

K2SO4(aq) ------------------> 2K^+ (aq) + SO4^2-(aq)

0.0574moles                       2*0.0574moles    0.0574moles

no of moles of K^+    = 2*0.0574moles    =0.1148 moles

no of moles of SO4^2-    = 0.0574moles

3.

volume of solution   = 250ml   = 0.25L

molarity of K2SO4   = no of moles K2SO4/volume of solution in L

                                = 0.0574/0.25   = 0.2296M

K2SO4(aq) ------------------> 2K^+ (aq) + SO4^2-(aq)

0.2296M                             2*0.2296M    0.2296M

concentration of K^+    = 2*0.2296M   = 0.4592M

concentration of SO4^2-    = 0.2296M